### v1

# V2P: From Background Suppression to Center Peaking for Robust GUI Grounding Task

Abstract
Precise localization of GUI elements is crucial for the development of GUI agents. Traditional methods rely on bounding box or center-point regression, neglecting spatial interaction uncertainty and visual-semantic hierarchies. Recent methods incorporate attention mechanisms but still face two key issues: (1) ignoring processing background regions causes attention drift from the desired area, and (2) uniform labeling fails to distinguish between center and edges of the target UI element, leading to click imprecision. Inspired by how humans visually process and interact with GUI elements, we propose the Valley-to-Peak (V2P) method to address these issues. To mitigate background distractions, V2P introduces a suppression attention mechanism that minimizes the model’s focus on irrelevant regions to highlight the intended region. For the issue of center-edge distinction, V2P applies a Fitts’ Law-inspired approach by modeling GUI interactions as 2D Gaussian heatmaps where the weight gradually decreases from the center towards the edges. The weight distribution follows a Gaussian function, with the variance determined by the target’s size. Consequently, V2P effectively isolates the target area and teaches the model to concentrate on the most essential point of the UI element. The model trained by V2P achieves the performance with 92.3% and 50.5% on two benchmarks ScreenSpot-v2 and ScreenSpot-Pro. Ablations further confirm each component’s contribution, highlighting V2P’s generalizability for precise GUI grounding tasks.
Figure 1: Performance of different baselines and our V2P-7B model on the dataset of ScreenSpot-v2 (left) and ScreenSpot-Pro (right). By training Qwen2.5-VL-7B using our Valley-to-Peak training strategy, our model achieves the best performance among all competitors.
1 Introduction
Recent advances in large language models (LLMs) and vision-language models (VLMs) have enabled agents to interpret natural language instructions and interact with graphical user interfaces (GUIs) across desktop, mobile, and web platforms. Central to this capability is GUI grounding—the alignment of language commands with semantically relevant UI elements and their spatial locations Cheng et al. (2024). This task bridges user intent and interface actions, supporting the development of intelligent, general-purpose agents for real-world human-computer interaction.
Early GUI grounding approaches framed the task as regression, outputting a bounding box or $(x,y)$ coordinate for a natural-language query Zhang et al. (2025); Qin et al. (2025). However, this “coordinate generation” method suffers weak spatial–semantic alignment Wu et al. (2025), treating coordinates like ordinary words without inherent spatial meaning. Moreover, point-wise regression contradicts the multi-point validity inherent in real interactions. Recent work addresses these issues by leveraging the model’s attention maps Wu et al. (2025). Instead of predicting coordinates, it extracts cross-modal attention weights linking instruction tokens to image patches, selecting the most attended patch as the click position. This approach offers dense spatial supervision and naturally tolerates multiple valid click regions, aligning better with human behavior.
Figure 2: Comparison of different strategies in the GUI grounding task. The green box marks the ground-truth bounding box, and the red box highlights the region where the model places the highest attention given the instruction and screenshot. The overlaid heatmap is colour-coded from cool (blue) to warm (red), with warmer colours indicating higher attention values.
However, empirical evidence reveals two persistent problems.
  1. 1.
    Background distraction: the current methods’ loss functions only require the model to increase attention on the target patches but do not constrain it to maintain low attention on the background patches. As a result, the model may assign high scores to background regions as well, leading to a "divergent" attention distribution. Due to the final softmax normalization, high-scoring background areas can absorb part of the normalized probability mass, which weakens or even shifts the attention peak on the target regions.
  2. 2.
    Centre-edge confusion: existing labels treat every pixel inside the bounding box equally, the model trained on such data is unable to differentiate between the center and edge regions of a UI element. As a result, it may allocate attention uniformly across both center and edge regions. This often leads to a prediction error where the click location deviates from the target center (i.e., an inaccurate click). Furthermore, when the labels of two UI elements partially overlap (e.g., a button overlapping with an input field), the uniform labeling scheme may cause the edge regions’ label values to accumulate excessively. This, in turn, can lead the model to overly focus on incorrect (and less relevant) areas while neglecting the truly important target center region.
This naturally raises the question: How can we guide the model’s visual attention to concentrate more precisely on the target UI element? Motivated by how humans first isolate a target (valley suppression) and then focus on its most clickable point (peak emphasis), we present Valley-to-Peak (V2P). V2P tackles the two issues synergistically:
Suppression Attention. To minimize distractions from irrelevant regions (such as background patches, icons, or text), we introduce an inverse attention regularization Li et al. (2018). This regularization penalizes high attention weights outside the target, creating low-attention "valleys" around non-target areas. By isolating the target while filtering out false positives, the network improves its focus on the true actionable UI elements in the interface.
Fitts-Gaussian Peak Modeling. Inspired by Fitts’ Law MacKenzie (1992); Fitts (1954), which formalizes human target acquisition performance as a function of target size and distance, we model the spatial distribution of interaction likelihood using a 2D Gaussian centered on the target element, with variance scaled proportionally to its spatial extent. This creates a probabilistic heatmap that is highest at the element’s core and decays smoothly towards the edges, naturally encoding the tolerance of human interaction. The resulting heatmap captures the spatial variability of interactions and steers the model toward a distinct attention peak at the most probable interaction location, thereby enhancing grounding precision.
Through joint optimization of these components, V2P reshapes the attention landscape, suppressing distractions to form valleys in irrelevant areas while emphasizing a sharp peak at the target’s actionable center. This balanced design significantly enhances grounding precision by aligning the model’s behavior with human interaction patterns.
We evaluate V2P on two public benchmarks: ScreenSpot-v2 Wu et al. (2024b) and the more challenging ScreenSpot-Pro Li et al. (2025). Our method achieves 92.3% on ScreenSpot-v2 and 50.5% on ScreenSpot-Pro, representing relative improvements of 3.5% and 23.7% respectively. Comprehensive ablation studies demonstrate that both Suppression Attention and Fitts-Gaussian Labeling consistently contribute to performance gains, highlighting the broad applicability of V2P to high-precision GUI grounding tasks.
2 Related Work
2.1 GUI-Agents
GUI agents have progressed from rudimentary random- or rule-based test tools to multimodal, LLM-driven systems that can follow natural-language instructions. Early efforts such as Monkey testing Wetzlmaier et al. (2016) and planning or script record-and-replay frameworks Memon et al. (2001); Steven et al. (2000) provided basic coverage but required hand-crafted rules or scripts. Machine-learning techniques later enabled more adaptive behaviour: Humanoid Li et al. (2020) and Deep GUI YazdaniBanafsheDaragh & Malek (2022) learned user-like action policies from screenshots, while widget detectors White et al. (2019) improved element recognition. Natural-language interfaces soon followed, e.g. FLIN Mazumder & Riva (2021) and RUSS Xu et al. (2021), and reinforcement learning environments like WoB Shi et al. (2017) and WebShop Yao et al. (2023) pushed web-scale interaction. The recent arrival of LLMs has unified perception, reasoning and control: WebAgent Gur et al. (2024) and WebGUM Furuta et al. (2024) achieve open-world browsing, AutoDroid Wen et al. (2024) and AppAgent Zhang et al. (2023) automate smartphones, and desktop agents such as UFO Zhang et al. (2024) demonstrate GPT-4-level capabilities; industrial systems (e.g. Claude 3.5 Sonnet and Operator) further attest to the practical traction of GUI agents.
2.2 GUI-Actor
GUI-Actor Wu et al. (2025) is the first VLM–based GUI agent that abandons explicit coordinate generation and instead performs coordinate-free grounding. A dedicated <ACTOR> token is appended to the vision–language encoder; its cross–modal attention to visual patch tokens is interpreted as a heat map from which the click location is taken. To cope with the fact that many pixels inside a UI element are equally valid, GUI-Actor adopts multi-patch supervision: every patch whose centre falls inside the ground-truth bounding box receives an identical positive label, whereas all remaining patches are treated as negatives. This design alleviates the granularity mismatch between patch-level features and pixel-level actions and yields a model that generalises better than previous coordinate-regression baselines across web, mobile and desktop benchmarks. GUI-Actor also introduces a lightweight grounding verifier that reranks the top-k attention peaks, further improving robustness.
Despite these advances, two issues remain. (1) Because the loss encourages high scores only on foreground patches and ignores the background, attention can drift toward visually salient but irrelevant areas. (2) Uniformly labelling all foreground patches fails to distinguish the centre of the element from its edges, which leads to imprecise clicks and error amplification when neighbouring elements overlap. Moreover, the additional verifier introduces a two-stage pipeline and extra inference cost, making the overall design less elegant.
Our Valley-to-Peak (V2P) framework inherits GUI-Actor’s coordinate-free backbone but resolves these issues with a single-stage solution: suppression attention explicitly down-weights backgrounds, and Fitts–Gaussian peak supervision highlights the element centre—eliminating the need for an external verifier while further boosting grounding accuracy.
3 Method
How can we guide a model’s attention to focus precisely on a target UI element while suppressing distractions? Inspired by human interaction patterns MacKenzie (1992); Fitts (1954), we present Valley-to-Peak (V2P), a framework that reshapes attention landscapes to better reflect real-world behavior through two key components:
  • Suppression Attention Valley Constraint: To reduce distractions, we introduce inverse attention regularization that penalizes high attention weights outside the target, forming low-attention “valleys” around irrelevant regions.
  • Fitts-Gaussian Peak Modeling: Uses a 2D Gaussian centered on the target to model interaction likelihood, creating a peak at the target’s core and guiding attention for precise grounding.
By jointly optimizing suppression and emphasis, V2P refines spatial focus on the actionable center while suppressing distractions. Also, unlike prior methods that treat UI elements rigidly, V2P employs size-adaptive Gaussian supervision to produce continuous, spatially graded attention maps that align with human patterns. The proposed training method significantly improves both the robustness and precision of UI localization, ensuring reliable performance under various challenges.
We describe the overall architecture in Section 3.1, then introduce suppression attention in Section 3.2 for reducing distractions. Finally, Section 3.3 details our Gaussian soft labeling scheme, which better aligns the model’s attention with natural human interaction patterns.
Figure 3: Valley-to-Peak training method (V2P). V2P jointly suppresses noise and enhances signals via two strategies: An inverse-attention penalty carves valleys in non-target areas, while size-adaptive Fitts-Gaussian peaks create sharp peaks at UI elements’ centers. This dual approach reshapes attention maps (rightmost example), enabling the model to quickly pinpoint interaction points in cluttered interfaces.
3.1 Model Architecture Overview
We build upon GUI-Actor Wu et al. (2025), a coordinate-free visual grounding framework that localizes GUI actions through attention rather than coordinate regression. Given a screenshot $I$ and an instruction $q$, the model introduces a special token <ACTOR> in the output sequence as a contextual anchor. The final-layer hidden state of <ACTOR>, denoted $h_{\texttt{<ACTOR>}}$, is used to compute action attention over image patch features $\{v_{1},\dots,v_{M}\}$ extracted by the vision encoder.
To enhance spatial coherence among visual patches, we apply a self-attention module over the patch features:
$\tilde{v}_{1},\dots,\tilde{v}_{M}=\text{SelfAttn}(v_{1},\dots,v_{M}),$ (1)
yielding contextualized representations. These are projected into a shared embedding space with $h_{\texttt{<ACTOR>}}$ via separate MLPs:
$z$ $=\text{MLP}_{T}(h_{\texttt{<ACTOR>}}),$ (2)
$z_{i}$ $=\text{MLP}_{V}(\tilde{v}_{i}),\quad i=1,\dots,M.$ (3)
Attention scores are then computed as:
$\alpha_{i}=\frac{z^{\top}z_{i}}{\sqrt{d}},\quad a_{i}=\frac{\exp(\alpha_{i})}{\sum_{j=1}^{M}\exp(\alpha_{j})},$ (4)
where $d$ is the embedding dimension. The resulting $\{a_{i}\}_{i=1}^{M}$ forms a normalized attention distribution over the $M$ image patches, representing the model’s belief about the target interaction location.
3.2 Suppression Attention Constraint for Distraction Mitigation
Attention maps in complex interfaces can suffer from attention leakage, where notable responses are mistakenly assigned to regions far from the target area, particularly in the presence of visually similar distracting patches. To address this issue and enhance spatial precision, we propose a Suppression Attention Constraint. This mechanism explicitly penalizes attention allocated to non-target regions, enforcing sparsity and improving the model’s ability to distinguish targets from surrounding distractions.
Let $\mathcal{G}\subset\{1,\dots,M\}$ denote the set of patch indices whose spatial support $R_{i}$ has empty intersection with the ground-truth bounding box $b$:
$\mathcal{G}=\left\{i\in\{1,\dots,M\}\mid R_{i}\cap b=\emptyset\right\}.$ (5)
We define the suppression attention loss as the total attention mass over these irrelevant regions:
$\mathcal{L}_{\text{Sup\_Attn}}=\sum_{i\in\mathcal{G}}a_{i}.$ (6)
This loss encourages the model to suppress attention on irrelevant regions, thereby reducing the impact of distracting elements in cluttered interfaces. By explicitly minimizing $\mathcal{L}_{\text{Sup\_Attn}}$, the model is incentivized to concentrate its focus on the target region, resulting in enhanced spatial precision and improved robustness.
3.3 Fitts-Gaussian Peak Modeling for Center-Focused Grounding
While the Suppression Attention Constraint encourages focus on target regions, overlapping UI elements can still lead to attention dispersion—particularly toward the boundaries of positively labeled components—resulting in ambiguous and spatially diffused attention maps.
To better align the model’s attention with human interaction behavior, we introduce Fitts-Gaussian Peak Modeling, a supervision strategy that explicitly encodes the expected attention peak over a target UI element. Our design is grounded in Fitts’ Law MacKenzie (1992); Fitts (1954), which models human pointing behavior: users naturally aim at the center of a target, and larger targets allow for greater spatial tolerance during interaction.
We model the ideal attention distribution as a 2D Gaussian density centered at the centroid of the ground-truth bounding box $b=[x_{1},y_{1},x_{2},y_{2}]$:
$\mu=(c_{x},c_{y})=\left(\frac{x_{1}+x_{2}}{2},\frac{y_{1}+y_{2}}{2}\right).$ (7)
To reflect the interaction tolerance associated with target size, we set the standard deviation of the Gaussian proportional to the element’s width and height:
$\sigma_{x}=\frac{w}{\sigma_{\text{factor}}},\quad\sigma_{y}=\frac{h}{\sigma_{\text{factor}}},$ (8)
where $w=x_{2}-x_{1}$, $h=y_{2}-y_{1}$, and $\sigma_{\text{factor}}$ is a hyperparameter controlling the concentration of the attention prior. This formulation ensures that larger elements—more tolerant to pointing errors—induce broader attention peaks, while smaller elements require sharper focus.
Given an input image partitioned into $M=H\times W$ non-overlapping patches of size $s\times s$, we compute the expected attention mass for each patch $i$, covering spatial region $R_{i}=[x^{i}_{\min},x^{i}_{\max}]\times[y^{i}_{\min},y^{i}_{\max}]$, by integrating the 2D Gaussian density over $R_{i}$:
$y_{i}=\int_{R_{i}}\mathcal{N}(x,y;\mu,\Sigma)dx\,dy,$ (9)
where $\Sigma=\mathrm{diag}(\sigma_{x}^{2},\sigma_{y}^{2})$. Thanks to axis-aligned separability, this integral decomposes efficiently into the product of two univariate cumulative distribution functions (CDFs):
$y_{i}=\left[\Phi(x^{i}_{\max};c_{x},\sigma_{x})-\Phi(x^{i}_{\min};c_{x},\sigma_{x})\right]\cdot\left[\Phi(y^{i}_{\max};c_{y},\sigma_{y})-\Phi(y^{i}_{\min};c_{y},\sigma_{y})\right],$ (10)
with $\Phi(\cdot\,;\mu,\sigma)$ denoting the CDF of a univariate normal distribution.
To supervise the model’s predicted attention distribution $\{a_{i}\}$, we adopt the action attention loss from GUI-Actor Wu et al. (2025), using the Kullback-Leibler (KL) divergence to measure the discrepancy between the target $p$ and prediction $a$:
$\mathcal{L}_{\text{Action\_Attn}}=\sum_{i=1}^{M}p_{i}\log\frac{p_{i}}{a_{i}},\quad p_{i}=\frac{y_{i}}{\sum_{j=1}^{M}y_{j}+\epsilon},\quad i=1,\ldots,M,$ (11)
where $\epsilon$ is a small constant for numerical stability.
Fitts-Gaussian Peak Modeling establishes a center-biased, size-aware attention prior that closely mimics human pointing behavior. By discouraging boundary leakage and promoting centralized attention in a graded, interaction-informed manner, it enhances localization precision and improves robustness in complex and cluttered UI layouts.
3.4 Valley-to-Peak Training
The overall training objective combines next-token prediction loss with action-focused attention losses:
$\mathcal{L}=\mathcal{L}_{\text{NTP}}+\lambda_{1}\mathcal{L}_{\text{Sup\_Attn}}+\lambda_{2}\mathcal{L}_{\text{Action\_Attn}},$ (12)
where $\mathcal{L}_{\text{Sup\_Attn}}$ suppresses attention outside the target region (Section 3.2), and $\mathcal{L}_{\text{Action\_Attn}}$ enforces alignment between predicted attention and a Gaussian-shaped target distribution (Section 3.3).
Minimizing the combined loss supports a Valley-to-Peak training paradigm: coarse suppression followed by fine-grained alignment. $\mathcal{L}_{\text{Sup\_Attn}}$ first suppresses distractions, guiding attention toward the target region. Then, $\mathcal{L}_{\text{Action\_Attn}}$ sharpens this focus by prioritizing the target’s center. This reduces misclicks and alleviates ambiguity caused by overlapping labels, ensuring precise and human-like attention alignment. The coarse-to-fine control enables robust interaction predictions, even in dense and visually complex UI environments.
4 Experiment
4.1 Experiment Setup
Implementation Details.
We use Qwen2.5-VL-7B-Instruct Bai et al. (2025) as the backbone VLM. Training is conducted on 32 NVIDIA H200 GPUs for 1 epoch. The key hyperparameters are a learning rate of 5e-6, a global batch size of 128, and $\sigma=1$ for the Fitts-Gaussian Peak Modeling.
Our data recipe is built from several public GUI datasets, the source data totaling approximately 1M screenshots. To ensure annotation quality, we apply Ominiparser Wan et al. (2024) to detect bounding boxes for all samples and filter those where the IoU between ground truth (GT) and parser-detected boxes is less than 0.3, as such cases likely contain annotation errors, this step improves the data consistency for training. More details about the implementation are provided in Appendix A.
Evaluation Benchmarks & Metric.
We conduct experiments to assess the performance of our methods on two widely adopted GUI visual grounding benchmarks: ScreenSpot-v2 Wu et al. (2024b), and ScreenSpot-Pro Li et al. (2025). Among these, ScreenSpot-Pro stands out due to its focus on higher-resolution user interfaces and increased domain variability, such as industrial applications and multi-window setups. This dataset serves as a valuable resource for evaluating generalization capabilities in realistic scenarios. To quantify performance, we utilize the Element Accuracy metric, which calculates the percentage of cases where the predicted click point resides within the target element’s ground-truth bounding box. Further details about these benchmarks are available in Appendix B
4.2 Main Result
Our proposed V2P-7B achieves outstanding performance on both the ScreenSpot-v2 and ScreenSpot-Pro benchmarks, delivering significant improvements over existing approaches. As shown in Tab. 1, V2P-7B achieves an average accuracy of 50.54% on the challenging ScreenSpot-Pro dataset, outperforming all GUI-specific models, including RL-based methods such as SE-GUI-7B (47.3%) and GUI-$\text{G}^{2}$-7B (47.5%). Notably, our model secures the highest scores in 6 out of 12 task categories, with exceptional performance in icon-based interactions (25.83% average), where precise spatial grounding is crucial. On ScreenSpot-v2 (Tab. 2), V2P-7B achieves an impressive average accuracy of 92.3%, with remarkable performance in Mobile-Text (98.1%) and Web-Text (95.4%) scenarios.
These advancements are driven by V2P’s dual optimization strategy: (1) the Suppression Attention mechanism effectively mitigates background distractions, while (2) Fitts-Gaussian Labeling addresses center-edge confusion, significantly improving localization in icon-based tasks. Remarkably, V2P obtains these results using supervised fine-tuning alone, avoiding the complexity of reinforcement learning pipelines employed by recent RL-based models. This streamlined training process not only simplifies implementation but also leaves substantial room for future performance enhancements through potential RL integration. The consistent gains across diverse UI platforms (mobile, desktop, web) and interaction types (text, icon) affirm V2P’s robust generalizability for real-world GUI grounding applications.
Model ScreenSpot-Pro Accuracy (%)
CAD Dev Creative Scientific Office OS Avg.
Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Avg.
Proprietary Models
GPT-4o 2.0 0.0 1.3 0.0 1.0 0.0 2.1 0.0 1.1 0.0 0.0 0.0 1.3 0.0 0.8
Claude Computer Use 14.5 3.7 22.0 3.9 25.9 3.4 33.9 15.8 30.1 16.3 11.0 4.5 23.4 7.1 17.1
General Open-source Models
Qwen2.5-VL-3B 9.1 7.3 22.1 1.4 26.8 2.1 38.2 7.3 33.9 15.1 10.3 1.1 23.6 3.8 16.1
Qwen2.5-VL-7B 16.8 1.6 46.8 4.1 35.9 7.7 49.3 7.3 52.5 20.8 37.4 6.7 38.9 7.1 26.8
GUI-specific Models (SFT)
SeeClick-9.6B 2.5 0.0 0.6 0.0 1.0 0.0 3.5 0.0 1.1 0.0 2.8 0.0 1.8 0.0 1.1
FOCUS-2B 7.6 3.1 22.8 1.7 23.7 1.7 25.0 7.1 23.2 7.7 17.8 2.5 19.8 3.9 13.3
CogAgent-18B 7.1 3.1 14.9 0.7 9.6 0.0 22.2 1.8 13.0 0.0 5.6 0.0 12.0 0.8 7.7
Aria-UI 7.6 1.6 16.2 0.0 23.7 2.1 27.1 6.4 20.3 1.9 4.7 0.0 17.1 2.0 11.3
OS-Atlas-7B 12.2 4.7 33.1 1.4 28.8 2.8 37.5 7.3 33.9 5.7 27.1 4.5 28.1 4.0 18.9
ShowUI-2B 2.5 0.0 16.9 1.4 9.1 0.0 13.2 7.3 15.3 7.5 10.3 2.2 10.8 2.6 7.7
UGround-7B 14.2 1.6 26.6 2.1 27.3 2.8 31.9 2.7 31.6 11.3 17.8 0.0 25.0 2.8 16.5
UGround-V1-7B 15.8 1.2 51.9 2.8 47.5 9.7 57.6 14.5 60.5 13.2 38.3 7.9 45.2 8.1 31.1
UI-TARS-2B 17.8 4.7 47.4 4.1 42.9 6.3 56.9 17.3 50.3 17.0 21.5 5.6 39.6 8.4 27.7
UI-TARS-7B 20.8 9.4 58.4 12.4 50.0 9.1 63.9 31.8 63.3 20.8 30.8 16.9 47.8 16.2 35.7
UI-TARS-72B 18.8 12.5 62.9 17.2 57.1 15.4 64.6 20.9 63.3 26.4 42.1 15.7 50.9 17.6 38.1
JEDI-3B 27.4 9.4 61.0 13.8 53.5 8.4 54.2 18.2 64.4 32.1 38.3 9.0 49.8 13.7 36.1
JEDI-7B 38.0 14.1 42.9 11.0 50.0 11.9 72.9 25.5 75.1 47.2 33.6 16.9 52.6 18.2 39.5
GUI-Actor-7B 44.6
GUI-specific Models (RL)
UI-R1-3B 11.2 6.3 22.7 4.1 27.3 3.5 42.4 11.8 32.2 11.3 13.1 4.5 24.9 6.4 17.8
UI-R1-E-3B 37.1 12.5 46.1 6.9 41.9 4.2 56.9 21.8 65.0 26.4 32.7 10.1 33.5
GUI-R1-3B 26.4 7.8 33.8 4.8 40.9 5.6 61.8 17.3 53.6 17.0 28.1 5.6
GUI-R1-7B 23.9 6.3 49.4 4.8 38.9 8.4 55.6 11.8 58.7 26.4 42.1 16.9
InfiGUI-R1-3B 33.0 14.1 51.3 12.4 44.9 7.0 58.3 20.0 65.5 28.3 43.9 12.4 49.1 14.1 35.7
GUI-G1-3B 39.6 9.4 50.7 10.3 36.6 11.9 61.8 30.0 67.2 32.1 23.5 10.6 49.5 16.8 37.1
SE-GUI-3B 38.1 12.5 55.8 7.6 47.0 4.9 61.8 16.4 59.9 24.5 40.2 12.4 50.4 11.8 35.9
SE-GUI-7B 51.3 42.2 68.2 19.3 57.6 9.1 75.0 28.2 78.5 43.4 49.5 25.8 63.5 21.0 47.3
GUI-$\text{G}^{2}$-7B 55.8 12.5 68.8 17.2 57.1 15.4 77.1 24.5 74.0 32.7 57.9 21.3 64.7 19.6 47.5
Ours
V2P-7B 58.38 12.50 67.53 24.83 62.63 16.08 73.61 33.64 75.71 43.40 56.07 32.58 65.81 25.83 50.54
Table 1: Comparison of Model Performance Across Task Categories in ScreenSpot-Pro. Bold text highlights the best results, while “–” represents missing values not reported in the original papers. Details of selected baselines are in Appendix C
Model ScreenSpot-v2 Accuracy (%)
Mobile-Text Mobile-Icon Desktop-Text Desktop-Icon Web-Text Web-Icon Avg.
Proprietary Models
Operator 47.3 41.5 90.2 80.3 92.8 84.3 70.5
GPT-4o + OmniParser-v2 95.5 74.6 92.3 60.9 88.0 59.6 80.7
General Open-source Models
Qwen2.5-VL-3B 93.4 73.5 88.1 58.6 88.0 71.4 80.9
Qwen2.5-VL-7B 97.6 87.2 90.2 74.2 93.2 81.3 88.8
GUI-specific Models (SFT)
SeeClick-9.6B 78.4 50.7 70.1 29.3 55.2 32.5 55.1
Magma-8B 62.8 53.4 80.0 57.9 67.5 47.3 61.5
OS-Atlas-4B 87.2 59.7 72.7 46.4 85.9 63.1 71.9
UI-TARS-2B 95.2 79.1 90.7 68.6 87.2 78.3 84.7
OS-Atlas-7B 95.2 75.8 90.7 63.6 90.6 77.3 84.1
Aguvis-7B 95.5 77.3 95.4 77.9 91.0 72.4 86.0
UGround-V1-7B 95.0 83.3 95.0 77.8 92.1 77.2 87.6
UI-TARS-72B 94.8 86.3 91.2 87.9 91.5 87.7 90.3
GUI-Actor-3B 97.6 83.4 96.9 83.6 94.0 85.7 91.0
UI-TARS-7B 96.9 89.1 95.4 85.0 93.6 85.2 91.6
GUI-Actor-7B 97.6 88.2 96.9 85.7 93.2 86.7 92.1
GUI-specific Models (RL)
SE-GUI-7B - - - - - - 90.3
LPO-8B - - - - - - 90.5
Ours
V2P-7B 98.1 88.0 96.1 89.7 95.4 84.4 92.3
Table 2: Comparison of Model Performance Across Task Categories in ScreenSpot-v2. Bold text highlights the best results, while “–” represents missing values not reported in the original papers.
4.3 Ablation Study
4.3.1 Ablation Study for V2P
The ablation study in Tab. 3 highlights the effectiveness of our V2P method. On the challenging ScreenSpot-Pro, removing Fitts-Gaussian Labeling results in a drop of 3.0% (50.5% to 47.5%), showing its importance in resolving center-edge confusion. Further removing Suppression Attention leads to an additional decline of 3.2%, underscoring its role in reducing background distractions. These results illustrate that first filtering out irrelevant regions to isolate target areas, followed by refining localization to pinpoint the center of UI elements with precision offering clear advantages in addressing complex scenarios.
On the simpler dataset ScreenSpot-v2, performance remains nearly unchanged after removing Fitts-Gaussian Labeling (92.3%), as the dataset’s straightforward tasks already allow high accuracy. And we suspect this is also because ScreenSpot-v2 primarily features simple UI layouts with minimal overlap, as Fitts-Gaussian Labeling is especially effective on more complex interfaces with overlapping UI elements, which guides the model to click precisely at the icon center, reducing misclicks on adjacent elements (we demonstrate it in Sec. 4.3.3). This indicates that while V2P provides limited gains on easier datasets like ScreenSpot-v2, its full potential shines in challenging real-world settings like ScreenSpot-Pro. However, removing both components results in a slight drop (91.9%), particularly in the Mobile-Icon scenario.
ScreenSpot-Pro
Model Dev Creative CAD Scientific Office OS Avg.
V2P-7B (Full) 46.8 43.1 47.1 56.3 68.3 45.4 50.5
   w/o Fitts-Gaussian Labeling 42.8 39.9 44.0 49.2 67.8 46.4 47.5 (-3.0 $\downarrow$)
   w/o Fitts-Gaussian & Suppression Attention 43.5 41.1 38.3 49.6 63.5 43.4 44.3 (-6.2 $\downarrow$)
ScreenSpot-v2
Model Mobile-Text Mobile-Icon Desktop-Text Desktop-Icon Web-Text Web-Icon Avg.
V2P-7B (Full) 98.1 88.0 96.11 89.7 95.4 84.4 92.3
   w/o Fitts-Gaussian Labeling 98.8 86.4 97.2 87.3 95.4 85.7 92.3 (-0.0)
   w/o Fitts-Gaussian & Suppression Attention 98.1 83.3 95.6 88.1 95.7 86.9 91.9 (-0.4 $\downarrow$)
Table 3: Ablation study on ScreenSpot-Pro and ScreenSpot-v2.
Figure 4: Ablation studies of Fitts-Gaussian Modeling (FGPM). (a) Effect of the Gaussian factor $\sigma$ on overall accuracy for ScreenSpot-v2 and ScreenSpot-Pro. (b) Size-wise performance on ScreenSpot-Pro.
4.3.2 Ablation Study for Gaussian Factor $\sigma$
We conducted ablation experiments to analyze the effect of different Gaussian factors $\sigma$ on model performance. As shown in Fig. 4(a), the model’s performance is strongly influenced by the choice of Gaussian factor $\sigma$. For both ScreenSpot-v2 and ScreenSpot-Pro, accuracy improves as $\sigma$ decreases. For example, on ScreenSpot-v2, the accuracy rises from 91.3% at $\sigma=6.0$ to 92.4% at $\sigma=0.5$, while ScreenSpot-Pro achieves its best result of 50.5% accuracy at $\sigma=1.0$.
We suspect that this is because that larger $\sigma$ values correspond to a broader Gaussian distribution, which tends to dilute the spatial focus and introduce noise into the attention maps. In comparison, smaller $\sigma$ produces sharper Gaussian peaks, allowing the model to localize UI elements with higher precision and resulting in more accurate click predictions. These results underscore that carefully balancing the Gaussian factor is crucial: excessively large values hinder localization, while moderate to smaller values significantly enhance spatial accuracy and overall model performance.
4.3.3 Ablation Study for Effectiveness of Fitts-Gaussian Peak Modeling
Traditional attention methods often yield overly broad regions, misaligning with small UI elements and producing points outside their boxes (Fig. 2(b)). Fitts-Gaussian Peak Modeling counters this by centering the attention, boosting accuracy on tiny elements. We ablate the design on the challenging ScreenSpot-Pro dataset Li et al. (2025), splitting its boxes into small, medium, and large and comparing results with/without our Fitts-Gaussian Peak Modeling (FGPM).
The results in Fig. 4(b) show that the proposed modeling yields substantial performance improvements for small UI elements by 7.4%, and also provides noticeable gains for medium elements (3.0%). However, for large UI elements, we observe a slight decrease in performance. We attribute this to the fact that our method’s centralized attention may sometimes neglect the full extent of the target, whereas the broader attention maps in prior methods might coincidentally localize such large elements better. Nevertheless, our approach still brings substantial overall improvement, especially in more challenging cases.
5 Conclusion
We propose Valley-to-Peak (V2P), a framework that addresses key challenges in leveraging vision-language model attentions for GUI grounding tasks (i.e., background distraction and center-edge confusion). By introducing Suppression Attention to reduce distractions and Fitts-Gaussian Labeling to guide the model toward actionable centers, V2P enhances the model’s ability to focus precisely on target elements, better simulating human interaction behavior. Experiments show V2P achieves modest gains on simpler datasets like ScreenSpot-v2 (92.3%, +3.5%) but excels in complex settings like ScreenSpot-Pro (50.5%, +23.7%). Ablation studies confirm the effectiveness of both components, highlighting V2P’s robustness and applicability to high-precision GUI grounding tasks.
Appendix A Training Details
A.1 Source Training Data
Following GUI-Actor Wu et al. (2025), we compile our training dataset from several publicly available, high-quality GUI datasets, with summary statistics provided in Table 4. To ensure fair evaluation, we also exclude any samples from Wave-UI that overlap with the test sets of downstream tasks.
Dataset # of Elements # of Screenshots Platform
Uground Web–Hybrid Gou et al. (2025a) 8M 775K Web
GUI-Env Chen et al. (2025) 262K 70K Web
GUI-Act Chen et al. (2025) 42K 13K Web
AndroidControl Li et al. (2024) 47K 47K Android
AMEX Chai et al. (2025) 1.2M 100K Android
Wave-UI 50K 7K Hybrid
Total 9.6M 1M
Table 4: Overview of training datasets used for GUI-Actor.
A.2 Training and Inference Details
During the training phase, we first freeze the backbone VLM parameters and train only the action head ($\sim 20$M parameters). In the second phase, we fine-tune the entire model using the filtered dataset with standard supervised learning. At inference, we follow deterministic generation with a temperature of 0 and adopt a confidence threshold of $\gamma=0.95$ for the ScreenSpot-Pro benchmark and $\gamma=0.8$ for ScreenSpot-v2 tasks.
Appendix B Benchmarks
Our evaluation centers on two sophisticated benchmarks for GUI visual grounding: ScreenSpot-v2 Wu et al. (2024b) and ScreenSpot-Pro Li et al. (2025).
ScreenSpot-v2 encompasses 1,272 carefully annotated instructions, each paired with corresponding target elements across diverse GUI environments, including mobile (Android and iOS), desktop (macOS and Windows), and web platforms. The dataset is designed to improve the quality and reliability of GUI visual grounding tasks, addressing key challenges such as eliminating ambiguities in natural language instructions and resolving annotation errors. By refining the alignment between textual descriptions and interface elements, ScreenSpot-v2 provides a robust and standardized benchmark for evaluating grounding models.
ScreenSpot-Pro, meanwhile, focuses on more demanding scenarios, especially those involving high-resolution professional applications. It contains 1,581 tasks annotated by domain experts across 23 specialized software applications, spanning three operating systems. This benchmark significantly broadens the scope of GUI visual grounding by introducing interfaces with industrial software and multi-window layouts, creating a larger domain gap compared to most pretraining data. With its increased complexity and domain diversity, ScreenSpot-Pro is an invaluable resource for assessing the generalization ability of models in realistic and challenging GUI environments.
Appendix C Baselines
C.1 Baselines for ScreenSpot-Pro
We establish comprehensive benchmarking across four categories of state-of-the-art GUI understanding systems:
  • Proprietary Systems: GPT-4o OpenAI (2024) (vision-language foundation model), Claude Computer Use Google (2024) (specialized GUI agent)
  • General-Purpose Open-Source: Qwen2.5-VL series Bai et al. (2025) (7B/72B parameter variants)
  • GUI-Specialized (SFT):
    • Medium-scale: SeeClick-9.6B Cheng et al. (2024), FOCUS-2B Zhang et al. (2024), OS-Atlas-7B Wu et al. (2024a)
    • Large-scale: CogAgent-18B Hong et al. (2024), Aria-UI Yang et al. (2025b), JEDI series Xie et al. (2025)
    • Domain-specific: ShowUI-2B Lin et al. (2024), Uground series Gou et al. (2025b), UI-TARS series Qin et al. (2025)
  • GUI-Specialized (RL):
    • R1-style: UI-R1 Lu et al. (2025), GUI-R1 Luo et al. (2025), InfiGUI-R1-3B Liu et al. (2025)
    • Gaussian-based: GUI-G1-3B Zhou et al. (2025), SE-GUI Yuan et al. (2025), GUI-$\text{G}^{2}$-7B Tang et al. (2025a)
C.2 Baselines for ScreenSpot-v2
We establish comprehensive benchmarking across four categories of state-of-the-art GUI understanding systems:
  • Proprietary Systems: Operator OpenAI (2023) (proprietary multimodal system), GPT-4o + OmniParser-v2 OpenAI (2024); Wan et al. (2024) (enhanced vision-language integration)
  • General-Purpose Open-Source: Qwen2.5-VL series Bai et al. (2025) (7B/72B parameter variants)
  • GUI-Specialized (SFT):
    • Medium-scale: SeeClick-9.6B Cheng et al. (2024), Magma-8B Yang et al. (2025a), OS-Atlas series Wu et al. (2024a)
    • Large-scale: UI-TARS series Qin et al. (2025), Uground series Gou et al. (2025b), GUI-Actor series Wu et al. (2025)
    • Domain-specific: Aguvis-7B Xu et al. (2025)
  • GUI-Specialized (RL):
    • SE-GUI-based: SE-GUI-7B Yuan et al. (2025), LPO-8B Tang et al. (2025b)